# Ballistic conductance with and without disorder in a boundary-driven XXZ spin chain

### v1

@@ Abstract
I Introduction
The Heisenberg chain and its descendants, especially the XXZ model, have long served as canonical models of magnetism and played a tremendously important role in the understanding of low-dimensional systems [1, 2, 3, 4, 5]. In particular, the quantum $S=1/2$ XXZ chain is a quintessential example of the connection between anomalous equilibrium hydrodynamics and integrability [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]; and even non-integrable spin chains can evince anomalous dynamics for remarkably long timescales [26, 27, 28, 29, 30, 31, 32].
Beyond the (near-)equilibrium paradigm, recent years have seen an increased interest in far-from-equilibrium phenomena [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] and the physics of open [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and driven (Floquet) systems [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85], where the transport phenomenology can differ markedly from the linear response regime.
In this vein, recent experimental work on Google’s noisy intermediate-scale quantum (NISQ) platform [86] has investigated the effect of edge driving on the $S=1/2$ XXZ chain, and yielded a detailed – in particular, fully spatially-resolved – spin transport phenomenology on a chain of $L=26$ sites, comprising non-equilibrium steady states (NESS) of ballistic, subdiffusive, and insulating nature (for easy-plane, isotropic, and easy-axis anisotropy, respectively), backed up by numerics as well as prior work on NESS [33].
@@ II Model
We consider a classical chain of unit-length spins ${\bm{S}_{i}\in S^{2}}$ with boundary-driven Floquet dynamics. A single time-step $t\mapsto t+\delta t$ consists of three parts. Between times $t$ and $t+\delta t/2$, odd bonds $(1,2)$, $(3,4)$, …, $(L-1,L)$ evolve, followed by the even bonds between times $t+\delta t/2$ and $t+\delta t$. Finally, the boundary spins are reset, $\bm{S}_{1}=-\bm{\hat{z}}$, $\bm{S}_{L}=+\bm{\hat{z}}$ (cf. iSWAP($\pi/2$) gates of the quantum experiment [86]).
@@
$\bm{S}_{1}$ $=-\bm{\hat{z}},\ \ \bm{S}_{2}=\chi\bm{\hat{x}}+\sqrt{1-\chi^{2}}\bm{\hat{z}},$
$\bm{S}_{3}$ $=...=\bm{S}_{L}=+\bm{\hat{z}},$ (4)
where, in what follows, we set $\chi=0.01$. This leads to a delay in the onset of the dynamics, and the injected current rises exponentially up to a time $t\approx 10^{2}\delta t$, Fig. 2(d). After this delay, the dynamics proceeds essentially in the same manner as in the quantum case, and, in particular, we show that the asymptotics are fundamentally equivalent.
III Clean spin transport
@@
III.1 Non-equilibrium steady states
@@
$\bm{S}_{i}(t)=\bm{S}_{i}(t+\delta t),\;\;\;\forall i.$ (5)
We consider first the limit of small time-steps $\delta t\to 0$ (or large driving frequency $\omega\to\infty$). This limit is well-defined, and corresponds to the continuous-time dynamics of the XXZ Hamiltonian (1), subject to the boundary conditions $\bm{S}_{1}=-\bm{\hat{z}}$, $\bm{S}_{L}=+\bm{\hat{z}}$. The steady-state condition becomes
$\dot{\bm{S}}_{i}(t)=0,\;\;\;\forall i,\;\;\;\omega\rightarrow\infty.$ (6)
@@ III.2 Transport regimes
III.3 Finite-frequency simulations
@@
We verify the predictions of the infinite-frequency calculation by performing numerical simulations of the dynamics at finite frequency. Starting from the initial state (4), we find the following (cf. Fig. 1): for easy-plane anisotropy ($\Delta<1$), a ballistic magnetisation front crosses the chain, and the NESS is attained at long times – the spin profile symmetrises between the $-\bm{\hat{z}}$ and $+\bm{\hat{z}}$ boundaries, oscillates around $S^{z}=0$ near the middle of the chain, and a spatially-uniform, non-zero spin current flows through the bulk.
@@
We plot the injected current for three representative anisotropies in Fig. 3(a). As noted in the discussion of the initial state, there is an initial exponential increase (from $j_{1}(0)\sim\chi^{2}$) as the system moves away from the unstable stationary state $(-\bm{\hat{z}},+\bm{\hat{z}},+\bm{\hat{z}},...)$, but after this delay we find the dynamical regimes predicted by the infinite-frequency calculation,
@@
III.4 Classical-quantum correspondence
@@
More concretely, the transport regimes, measured by the decay of the injected current $j_{1}(t)$, are precisely equivalent in both the classical model and the quantum experiments, and are in accordance with the predictions of the infinite-frequency analysis of the steady-states in §III.1.
IV Disordered spin transport
@@
Having shown that the clean classical model has the same transport regimes as the quantum model [86], we next investigate how the phenomenology is affected by the introduction of random couplings. We will primarily consider the easy-plane regime, and, in §IV.4, the isotropic point. The easy-axis case is not particularly interesting: there is no transport in the clean case, and, unsurprisingly, there is also no transport in the disordered case.
@@ IV.1 Transport observables
We deal with (i) by considering the full probability distribution of steady-state currents $\rho(j_{\infty})$ and not just the disorder-average $\overline{j_{\infty}}$. We will return to (ii) when we perform finite-frequency simulations in §IV.3.
IV.2 Stability of the ballistic regime
Figure 6: Ballistic conductance in the disordered chain. Histograms of the distribution of steady-state currents, $\rho(j_{\infty}(L))$ for $\Delta=0.8$ and $W=0.2$ (a), $W=0.5$ (b), and $W=0.8$ (c). Each histogram is constructed from the final values $j_{L-1}(t_{f})$ for $t_{f}=10^{7}\delta t$, $\delta t=\pi/2$, over $10^{4}$ realisations of disorder. We find that $\rho(j_{\infty})$ is independent of system size, at least up to $L=1024$, and that, as predicted by the infinite-frequency analysis, the steady-state current distribution has a nonzero lower bound. Note that for $W=0.8$, due to the increased relaxation times (cf. Fig 7), we have, only for this figure, started from the initial state $\bm{S}_{2}=...=\bm{S}_{L-1}=\bm{\hat{x}}$ in order that the steady states are reached before the end of the simulation.
@@ IV.2 Stability of the ballistic regime
We conjecture that the minimum conductance $G(L)$ (as a function of the disorder realisation) is obtained from the couplings $J_{2}=...J_{L-2}=1-W$, and that this, indeed, corresponds to a finite, length-independent lower bound on $G(L)$. We show this minimum conductance as a function of the disorder strength $W$, for various anisotropies $\Delta<1$, in Fig. 4(a).
@@
IV.3 Finite-frequency simulations with disorder
@@
We should address, however, the question of the timescale on which the length-independent steady-state distribution is set up. Indeed, this timescale need not (and, we will show, does not) have the same length-scaling as the steady-state current: the arguments of §IV.2 imply that the steady-state current is primarily determined by the bonds near the edge of the chain; but to attain the NESS from the initial state (4) the magnetisation must propagate across the full length of the chain, which clearly involves every bond.
@@
It is here that the disordered nature of the chain becomes visible – the distribution of crossing times $t^{*}$ exhibits diffusive, not ballistic, behaviour, on the longest length-scales we can simulate. We show in Fig. 7 that the crossing time scales as $t^{*}\sim L^{2}$ (i.e., diffusively); though we show in Fig. 7(a) that it can appear to be superdiffusive or even ballistic up to surprisingly large system sizes $L\gtrsim 10^{3}$ (depending on $W$), which would be relevant to experiments on near-term quantum devices.
IV.4 Subdiffusion at strong disorder
Finally, let us briefly address what happens when the disorder strength reaches $W=1$. The arguments of §IV.2 no longer apply, and, as shown in Fig. 5(a), we find a subdiffusive scaling of the disorder average of the injected current, $\overline{j_{1}(t)}\sim t^{-2/3}$. We discuss a possible explanation for this observation in App. D.
@@ V Conclusions
Acknowledgements.
We thank Dima Abanin, Pieter Claeys, Ferdinand Evers, Sarang Gopalakrishnan, David Huse and Xiao Mi for many useful discussions. This work was in part supported by the Deutsche Forschungsgemeinschaft under grants Research Unit FOR 5522 (project-id 499180199) and the cluster of excellence ct.qmat (EXC 2147, project-id 390858490).
@@ Appendix A Finite-frequency dynamics
In this appendix we derive the relationship between the values of the spins in the steady-state and the conductance of the chain. We first note that any steady-state current must be spatially uniform, so it suffices to consider only the final two spins $\bm{S}_{L-1}$ and $\bm{S}_{L}$. Now, at the start of a given timestep, $\bm{S}_{L}(t)=+\bm{\hat{z}}$, and, at the end of the timestep $t+\delta t$, $\bm{S}_{L}$ is reset to $+\bm{\hat{z}}$. The magnetisation extracted through the boundary is, therefore,
$j_{L-1}(t)=1-S^{z}_{L}(t+\delta t/2)$ (11)
(where we have assumed that $L$ is even such that the bond $(L-1,L)$ evolves over the first half of the timestep, but this detail is not important). At the isotropic point $\Delta=1$, the dynamics of a single bond is particularly straightforward: both spins precess around the total conserved value $\bm{S}_{L-1}+\bm{S}_{L}$. Denoting $\bm{S}_{L-1}(t)$ by $z_{L-1}$, one thus obtains
$j_{L-1}=\frac{1}{2}(1-z_{L-1})(1-\cos(\delta t/2)).$ (12)
By assumption, this is the steady-state current, and thus ${j_{\infty}(L)\sim G(L)\sim 1-z_{L-1}}$. We note that this calculation does not depend on whether we are using the infinite-frequency NESS, or the finite-frequency NESS satisfying the stroboscopic condition (5). The proportionality, ${j_{\infty}(L)\sim 1-z_{L-1}}$, continues to hold away from the isotropic point.
@@ Appendix B Solution of the consistency equations
$z_{n}=\sin\left(\frac{\pi}{L}\left(n-\frac{L+1}{2}\right)\right),$ (13)
@@
$z_{L-1}(L)=1-\frac{\pi^{2}}{8}L^{-2}+\frac{\pi^{4}}{384}L^{-4}+\dots,$ (14)
@@
$z_{n}\sim\tanh\left[\left(n-\frac{L+1}{2}\right)\mathrm{arccosh}(\Delta)\right],\;\;L\to\infty$ (16)
@@ Appendix D Subdiffusion at strong disorder
We turn our attention now to the transition away from the ballistic regime at strong disorder: the arguments of §IV.2 do not apply if there is no nonzero lower bound to the coupling strengths, $J_{i}\geq J_{\mathrm{min}}>0$, and, as shown in Fig. 5(a), the exponent jumps from ballistic ($\alpha=1$) to subdiffusive ($\alpha=1/3$) behaviour at $W=1$.
@@
The previous analysis, however, has neglected the fact that, for any NESS, a very small $J_{i}$ in the bulk limits the total current.
Figure 9: Time-dependence of the disorder-averaged injected current $\overline{j_{1}(t)}$, for the power-law distributions of disorder, Eq. (17). The power-law asymptotes use the subdiffusive exponent $\alpha(\eta)$ predicted by Eq. (19). Again, $10^{4}$ realisations of disorder have been averaged over. $L=1024$.
For $W<1$, we could reasonably neglect this detail as our primary interest lies only in the transport exponent, and the lower bound to $J_{i}$ implied that each bond could support some minimum current.
@@
$\overline{\min\limits_{i=1,...,L}J_{i}}\sim L^{-\frac{1}{1-\eta}}.$ (18)
Assuming, since this is a global bottleneck, that it is directly proportional to the conductivity (not the conductance), we obtain
$G(L)\sim L^{-\frac{1}{1-\eta}-1}\;\;\;\Rightarrow\;\;\;\alpha=\frac{1}{2+\frac{1}{1-\eta}},$ (19)

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
The Heisenberg chain and its descendants, especially the XXZ model, have long served as canonical models of magnetism and played a tremendously important role in the understanding of low-dimensional systems [1, 2, 3, 4, 5]. In particular, the quantum $S=\frac{1}{2}$ XXZ chain is a quintessential example of the connection between anomalous equilibrium hydrodynamics and integrability [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]; and even non-integrable spin chains can evince anomalous dynamics for remarkably long timescales [26, 27, 28, 29, 30, 31, 32].
Beyond the (near-)equilibrium paradigm, recent years have seen an increased interest in far-from-equilibrium phenomena [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] and the physics of open [53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] and driven (Floquet) systems [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86], where the transport phenomenology can differ markedly from the linear response regime.
In this vein, recent experimental work on Google’s noisy intermediate-scale quantum (NISQ) platform [87] has investigated the effect of edge driving on the $S=\frac{1}{2}$ XXZ chain, and yielded a detailed – in particular, fully spatially-resolved – spin transport phenomenology on a chain of $L=26$ sites, comprising non-equilibrium steady states (NESS) of ballistic, subdiffusive, and insulating nature (for easy-plane, isotropic, and easy-axis anisotropy, respectively), backed up by numerics as well as prior work on NESS [33].
@@ 2 Model
We consider a classical chain of unit-length spins ${\boldsymbol{S}_{i}\in S^{2}}$ with boundary-driven Floquet dynamics. A single time-step $t\mapsto t+\delta t$ consists of three parts. Between times $t$ and $t+\delta t/2$, odd bonds $(1,2)$, $(3,4)$, …, $(L-1,L)$ evolve, followed by the even bonds between times $t+\delta t/2$ and $t+\delta t$. Finally, the boundary spins are reset, $\boldsymbol{S}_{1}=-\boldsymbol{\hat{z}}$, $\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}}$ (cf. iSWAP($\pi/2$) gates of the quantum experiment [87]).
@@
$\boldsymbol{S}_{1}$ $=-\boldsymbol{\hat{z}},\ \ \boldsymbol{S}_{2}=\chi\boldsymbol{\hat{x}}+\sqrt{1-\chi^{2}}\boldsymbol{\hat{z}},$
$\boldsymbol{S}_{3}$ $=...=\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}},$ (4)
where, in what follows, we set $\chi=0.01$. This leads to a delay in the onset of the dynamics, and the injected current rises exponentially up to a time $t\approx 10^{2}\delta t$, Fig. 3(a). After this delay, the dynamics proceeds essentially in the same manner as in the quantum case, and, in particular, we show that the asymptotics are fundamentally equivalent.
3 Clean spin transport
@@
3.1 Non-equilibrium steady states
@@
$\boldsymbol{S}_{i}(t)=\boldsymbol{S}_{i}(t+\delta t),\;\;\;\forall i.$ (5)
We consider first the limit of small time-steps $\delta t\to 0$ (or large driving frequency $\omega\to\infty$). This limit is well-defined, and corresponds to the continuous-time dynamics of the XXZ Hamiltonian (1), subject to the boundary conditions $\boldsymbol{S}_{1}=-\boldsymbol{\hat{z}}$, $\boldsymbol{S}_{L}=+\boldsymbol{\hat{z}}$. The steady-state condition becomes
$\dot{\boldsymbol{S}}_{i}(t)=0,\;\;\;\forall i,\;\;\;\omega\rightarrow\infty.$ (6)
@@ 3.2 Transport regimes
3.3 Finite-frequency simulations
@@
We verify the predictions of the infinite-frequency calculation by performing numerical simulations of the dynamics at finite frequency. Starting from the initial state (4), we find the following (cf. Fig. 1): for easy-plane anisotropy ($\Delta<1$), a ballistic magnetisation front crosses the chain, and the NESS is attained at long times – the spin profile symmetrises between the $-\boldsymbol{\hat{z}}$ and $+\boldsymbol{\hat{z}}$ boundaries, oscillates around $S^{z}=0$ near the middle of the chain, and a spatially-uniform, non-zero spin current flows through the bulk.
@@
We plot the injected current for three representative anisotropies in Fig. 3(a). As noted in the discussion of the initial state, there is an initial exponential increase (from $j_{1}(0)\sim\chi^{2}$) as the system moves away from the unstable stationary state $(-\boldsymbol{\hat{z}},+\boldsymbol{\hat{z}},+\boldsymbol{\hat{z}},...)$, but after this delay we find the dynamical regimes predicted by the infinite-frequency calculation,
@@
3.4 Classical-quantum correspondence
@@
More concretely, the transport regimes, measured by the decay of the injected current $j_{1}(t)$, are precisely equivalent in both the classical model and the quantum experiments, and are in accordance with the predictions of the infinite-frequency analysis of the steady-states in §3.1.
4 Disordered spin transport
@@
Having shown that the clean classical model has the same transport regimes as the quantum model [87], we next investigate how the phenomenology is affected by the introduction of random couplings. We will primarily consider the easy-plane regime, and, in §4.4, the isotropic point. The easy-axis case is not particularly interesting: there is no transport in the clean case, and, unsurprisingly, there is also no transport in the disordered case.
@@ 4.1 Transport observables
We deal with (i) by considering the full probability distribution of steady-state currents $\rho(j_{\infty})$ and not just the disorder-average $\overline{j_{\infty}}$. We will return to (ii) when we perform finite-frequency simulations in §4.3.
4.2 Stability of the ballistic regime
Figure 6: Ballistic conductance in the disordered chain. Histograms of the distribution of steady-state currents, $\rho(j_{\infty}(L))$ for $\Delta=0.8$ and $W=0.2$ (a), $W=0.5$ (b), and $W=0.8$ (c). Each histogram is constructed from the final values $j_{L-1}(t_{f})$ for $t_{f}=10^{6}\delta t$, $\delta t=\pi/2$, over $10^{4}$ realisations of disorder. We find that $\rho(j_{\infty})$ is independent of system size, at least up to $L=1024$, and that, as predicted by the infinite-frequency analysis, the steady-state current distribution has a nonzero lower bound. Note that for $W=0.8$, due to the increased relaxation times (cf. Fig 7), we have, only for this figure, started from the initial state $\boldsymbol{S}_{2}=...=\boldsymbol{S}_{L-1}=\boldsymbol{\hat{x}}$ in order that the steady states are reached before the end of the simulation.
@@ 4.2 Stability of the ballistic regime
We conjecture that the minimum conductance $G(L)$ (as a function of the disorder realisation) is obtained from the couplings $J_{2}=...=J_{L-2}=1-W$, and that this, indeed, corresponds to a finite, length-independent lower bound on $G(L)$. We show this minimum conductance as a function of the disorder strength $W$, for various anisotropies $\Delta<1$, in Fig. 4(a).
@@
4.3 Finite-frequency simulations with disorder
@@
We should address, however, the question of the timescale on which the length-independent steady-state distribution is set up. Indeed, this timescale need not (and, we will show, does not) have the same length-scaling as the steady-state current: the arguments of §4.2 imply that the steady-state current is primarily determined by the bonds near the edge of the chain; but to attain the NESS from the initial state (4) the magnetisation must propagate across the full length of the chain, which clearly involves every bond.
@@
[…]
4.4 Subdiffusion at strong disorder
Finally, let us briefly address what happens when the disorder strength reaches $W=1$. The arguments of §4.2 no longer apply, and, as shown in Fig. 5(a), we find a subdiffusive scaling of the disorder average of the injected current, $\overline{j_{1}(t)}\sim t^{-2/3}$. We discuss a possible explanation for this observation in App. D.
@@ 5 Conclusions
Acknowledgements
We thank Dima Abanin, Pieter Claeys, Ferdinand Evers, Sarang Gopalakrishnan, David Huse and Xiao Mi for many useful discussions.
Funding information
This work was in part supported by the Deutsche Forschungsgemeinschaft under grants Research Unit FOR 5522 (project-id 499180199) and the cluster of excellence ct.qmat (EXC 2147, project-id 390858490).
@@ Appendix A Finite-frequency dynamics
In this appendix we derive the relationship between the values of the spins in the steady-state and the conductance of the chain. We first note that any steady-state current must be spatially uniform, so it suffices to consider only the final two spins $\boldsymbol{S}_{L-1}$ and $\boldsymbol{S}_{L}$. Now, at the start of a given timestep, $\boldsymbol{S}_{L}(t)=+\boldsymbol{\hat{z}}$, and, at the end of the timestep $t+\delta t$, $\boldsymbol{S}_{L}$ is reset to $+\boldsymbol{\hat{z}}$. The magnetisation extracted through the boundary is, therefore,
$j_{L-1}(t)=1-S^{z}_{L}(t+\delta t/2)$ (A.1)
(where we have assumed that $L$ is even such that the bond $(L-1,L)$ evolves over the first half of the timestep, but this detail is not important). At the isotropic point $\Delta=1$, the dynamics of a single bond is particularly straightforward: both spins precess around the total conserved value $\boldsymbol{S}_{L-1}+\boldsymbol{S}_{L}$. Denoting $\boldsymbol{S}_{L-1}(t)$ by $z_{L-1}$, one thus obtains
$j_{L-1}=\frac{1}{2}(1-z_{L-1})(1-\cos(\delta t/2)).$ (A.2)
By assumption, this is the steady-state current, and thus ${j_{\infty}(L)\sim G(L)\sim 1-z_{L-1}}$. We note that this calculation does not depend on whether we are using the infinite-frequency NESS, or the finite-frequency NESS satisfying the stroboscopic condition (5). The proportionality, ${j_{\infty}(L)\sim 1-z_{L-1}}$, continues to hold away from the isotropic point, cf. Fig. 3(b).
@@ Appendix B Solution of the consistency equations
$z_{n}=\sin\left(\frac{\pi}{L}\left(n-\frac{L+1}{2}\right)\right),$ (B.1)
@@
$z_{L-1}(L)=1-\frac{\pi^{2}}{8}L^{-2}+\frac{\pi^{4}}{384}L^{-4}+\dots,$ (B.2)
@@
$z_{n}\sim\tanh\left[\left(n-\frac{L+1}{2}\right)\mathrm{arccosh}(\Delta)\right],\;\;L\to\infty$ (B.4)
@@ Appendix D Subdiffusion at strong disorder
We turn our attention now to the transition away from the ballistic regime at strong disorder: the arguments of §4.2 do not apply if there is no nonzero lower bound to the coupling strengths, $J_{i}\geq J_{\mathrm{min}}>0$, and, as shown in Fig. 5(a), the exponent jumps from ballistic ($\alpha=1$) to subdiffusive ($\alpha=1/3$) behaviour at $W=1$.
@@
The previous analysis, however, has neglected the fact that, for any NESS, a very small $J_{i}$ in the bulk limits the total current. For $W<1$, we could reasonably neglect this detail as our primary interest lies in the transport exponent, and the lower bound to $J_{i}$ implied that each bond could support some minimum current.
Figure 9: Time-dependence of the disorder-averaged injected current $\overline{j_{1}(t)}$, for the power-law distributions of disorder, Eq. (D.1). The power-law asymptotes use the subdiffusive exponent $\alpha(\eta)$ predicted by Eq. (D.3). Again, $10^{4}$ realisations of disorder have been averaged over. $L=1024$.
@@
$\overline{\min\limits_{i=1,...,L}J_{i}}\sim L^{-\frac{1}{1-\eta}}.$ (D.2)
Assuming, since this is a global bottleneck, that this value is directly proportional to the conductivity (not the conductance), we obtain
$G(L)\sim L^{-\frac{1}{1-\eta}-1}\;\;\;\Rightarrow\;\;\;\alpha=\frac{1}{2+\frac{1}{1-\eta}},$ (D.3)